# ShaRP: SHAllow-LayeR Pruning for Video Large Language Models Acceleration

### v1

@@ 5.1 Experimental Settings
Baselines. We compare our method with six representative training-free token pruning approaches. FastV [6] and PDrop [7] perform attention-guided pruning starting from the 2nd and 4th decoder layers, respectively. Feather [8] removes positional embeddings at the 2nd layer and applies uniform token sampling before pruning. FrameFusion [24] conducts progressive, layer-wise pruning based on token similarity and selects important tokens according to attention scores. To verify the compatibility of our strategy, we further integrate it into two representative pruning methods, VisionZip [10] and HoliTom [19]. Specifically, VisionZip performs outer-LMM frame-level pruning by identifying important tokens with a vision encoder, on top of which we apply our inner-LMM compression for further token reduction. HoliTom [19] adopts a two-stage pruning framework, where redundant tokens are progressively removed through outer- and inner-LMM pruning. We replace its inner-LMM pruning stage with our proposed strategy. For layer-wise compression methods, FrameFusion [24], we compute an equivalent retention ratio based on the average proportion of retained tokens across all LLM layers to match overall computational cost. Unless otherwise specified, we follow the official configurations of both backbones.
Implementation Details. We implement ShaRP on two representative multimodal LLMs, i.e. LLaVA-OneVision-7B [3] and Qwen2-VL-7B [4], which embody distinct architectural designs to ensure the robustness of our approach and we apply pruning at the first decoder layer. All experiments are conducted on NVIDIA A40 GPUs using the LMMs-Eval framework [65]. The hyperparameters are empirically determined and fixed to , , for all experiments. See more details in the supplementary materials.
@@ 5.2 Main Results
Enhanced Overall Performance. For LLaVA-OneVision [3], tokens from 32 sampled frames are evaluated under retention ratios about , with FLOPs matched across methods for fair comparison. Tab. 1 reports average scores and accuracy under different and TFLOPs. FastV [6] achieves decent performance at higher retention ratios (e.g., ), yet its accuracy declines sharply under aggressive pruning (e.g., ), indicating that raw attention scores without correction are unreliable when few tokens remain. Feather [8] alleviates positional bias by removing local positional encoding (at a certain layer) and uniform sampling. It shows limited performance under low retention ratios as many tokens retained are irrelevant to the query, achieving only 95.4% of the vanilla model’s performance even at 25% retention. PDrop [7] employs a progressive, layered pruning method. To achieve comparable FLOPs, this strategy necessitates aggressive pruning in deeper layers, which leads to very low accuracy even if the shallow layers retain the complete set of tokens. In contrast, ShaRP preserves up to 99% of vanilla accuracy even with aggressive shallow-layer pruning (72%). Its advantage is particularly pronounced on long-form video benchmarks such as LongVideoBench [63] and MLVU [64], where it surpasses the second-best method by 2.8% and 3.1% at 10% retention, respectively. These gains arise from ShaRP’s ability to aggregate information across temporally distant frames. Across all retention ratios, ShaRP consistently achieves competitive performance, demonstrating its robustness and adaptability.
Cross-Backbone Generalization. To evaluate the cross-model generalization of our approach, we extend experiments to Qwen2-VL [4], as detailed in Tab. 2. For this model, 32 sampled frames generate a token sequence as input, and we test with retention ratios across MVBench [61, 62], LongVideoBench [63], and VideoMME [52]. Unlike the LLaVA series, Qwen2-VL processes video frames using 3D convolutions and uses an MLP to aggregate tokens. For fair comparison under reasonable memory constraints, all methods are evaluated at the same input resolution. As shown in Tab 2, ShaRP consistently outperforms all baselines across the evaluated retention ratios, demonstrating its robustness across different backbones.
@@
Inference Efficiency. Tab. 4 presents the inference efficiency of ShaRP on the LLaVA-OneVision [3] benchmark. Our approach maintains strong performance while significantly reducing computational complexity. For instance, at a retention rate of 22.8%, TFLOPS decreases to 9.31, cutting pre-filling latency from 1181.6 ms to 538.5 ms (about 2.2 speedup). With a more aggressive rate of 14.7%, latency is further reduced to 393.4 ms, achieving approximately 3.0 acceleration. We further analyze the pre-filling time distribution across stages: debiasing + local masking + pruning (denoted as "Attn." in Tab. 4), token pre-filtering ("Fltr."), deduplication ("Dedpl."), post-filling ("Fil."), and the standard VLLM forward pass ("Fw."). Results show that debiasing and masking incur minimal overhead, whereas token deduplication accounts for nearly half of the total time, which highlights a key target for future optimization.
@@
MVBench ↑ |
MLVU ↑ | VideoMME ↑ |
| VisionZip [10] | 12.8% | 4.3 | 55.6 | 51.6 | 60.9 | 55.4 | 55.9 | ||||||||
[…]
| HoliTom [19] | 15%/7.5% | 4.3 | 58.4 | 55.7 | 62.1 | 57.2 | 58.4 | ||||||||
[…]
| TFLOPs | Prefilling Time (ms) ↓ | Average ↑ | |||||
| Attn. ↓ | Fltr. ↓ | Dedpl. ↓ | Fil. ↓ | LLM Fw. ↓ | Total ↓ | ||
| 40.8 (100%) | – | – | – | – | 1181.6 | 1181.6 (1.0) | 59.0 (100%) |
| 9.31 (22.8%) | 26.3 | 13.1 | 234.7 | 54.0 | 212.6 | 538.5 (2.2) | 58.4 (98.9%) |
| 6.03 (14.7%) | 26.3 | 12.1 | 143.6 | 52.0 | 159.4 | 393.4 (3.0) | 57.4 (97.2%) |
@@ 5.3 Ablation Study
Ablation and Synergy Analysis. We conduct ablation experiments on our three modules—Segment-Aware Causal Masking (SegM, Sec. 4.1), Positional Bias Calibration (PosC, Sec. 4.2), and Register Token Deduplication (RegD, Sec. 4.3)—under a 23.5% token retention ratio, with results in Tab. 5. PosC alone improves the average score from 55.8 to 57.4 by enhancing focus on critical visual regions. Adding SegM (PosC+SegM) further increases performance to 57.6, mitigating attention collapse and stabilizing temporal-contextual aggregation. Combining PosC with RegD raises the score to 58.0, demonstrating RegD’s effectiveness in compressing redundant tokens and enriching structural representation. Even without PosC, SegM+RegD reaches 57.6, confirming their independent contributions. Integrating all three modules achieves the best score of 58.4, validating the complementary benefits of each component
| Method | MVBench ↑ |
MLVU ↑ | VideoMME ↑ |
| Original | 54.4 | 53.0 | 61.0 | 54.7 | 55.8 | ||||
| PosC | 56.6 | 55.3 | 60.8 | 57.0 | 57.4 | ||||
| PosC+SegM | 56.6 | 55.5 | 60.9 | 57.2 | 57.6 | ||||
| PosC+RegD | 56.6 | 56.0 | 62.5 | 57.0 | 58.0 | ||||
| SegM+RegD | 56.5 | 55.0 | 62.2 | 56.7 | 57.6 | ||||
| PosC+SegM+RegD (ShaRP) | 57.1 | 56.1 | 62.4 | 58.0 | 58.4 |

### v2

@@ 5.1 Experimental Settings
Baselines. We compare ShaRP with eight representative training-free token pruning approaches. For image compression, FastV [7] and PDrop [8] perform attention-guided pruning starting from the 2nd and 4th decoder layers, respectively. Feather [10] removes positional embeddings at the 2nd layer and applies uniform token sampling before pruning. VisionZip [12] performs outer-LLM frame-level pruning using a vision encoder, upon which we apply ShaRP as an inner-LLM compression stage for additional reduction. For video token compression, FrameFusion [30] conducts progressive layer-wise pruning based on token similarity and attention importance, we compute an equivalent retention ratio based on the average proportion of retained tokens to ensure comparable computational cost. DyCoke [15] performs temporal merging prior to LLM processing, while STTM [20] reduces redundancy through spatiotemporal token merging. HoliTom [22] adopts a two-stage design with outer- and inner-LLM pruning; we replace its inner pruning stage with ShaRP to evaluate compatibility. Unless otherwise specified, we follow the official backbone configurations.
Implementation Details. We implement ShaRP on four representative multimodal LLMs, i.e., LLaVA-OneVision-7B [2], LLaVA-Video-7B [3], Qwen2-VL-7B [4], and Qwen2.5-VL-7B [5], which embody distinct architectural designs to ensure the robustness of our approach and we apply pruning at the first decoder layer. All experiments are conducted on NVIDIA A40 GPUs using the LMMs-Eval framework [71]. The hyperparameters are empirically determined and fixed to , , for all experiments. See more details in the supplementary materials.
@@ 5.2 Main Results
Enhanced Overall Performance. For LLaVA-OneVision [2], tokens from 32 sampled frames are evaluated under retention ratios about , with FLOPs matched across methods for fair comparison. Tab. 1 reports average scores and accuracy under different and TFLOPs. FastV [7] achieves decent performance at higher retention ratios (e.g., ), yet its accuracy declines sharply under aggressive pruning (e.g., ), indicating that raw attention scores are unreliable when few tokens remain. Feather [10] alleviates positional bias by removing local positional encoding (at a certain layer) and uniform sampling. It shows limited performance under low retention ratios as many tokens retained are irrelevant to the query, achieving only 95.4% of the vanilla model’s performance even at 25% retention. PDrop [8] employs a progressive, layered pruning method. To achieve comparable FLOPs, this strategy necessitates aggressive pruning in deeper layers, which leads to very low accuracy even if the shallow layers retain the complete set of tokens. DyCoke [15] and STTM [20] reduce redundancy through temporal or spatiotemporal merging; however, they exhibit notable performance degradation even at moderate compression levels. For example, at approximately retention, their average performance drops to 88.7% and 91.3%, respectively. In contrast, ShaRP preserves up to 99.3% of vanilla accuracy even with aggressive shallow-layer pruning (72%). Its advantage is particularly pronounced on long-form video benchmarks such as LongVideoBench [69] and MLVU [70], where it surpasses the second-best method by 2.3% and 3.7% at 6.03 TFLOPs, respectively. These gains arise from ShaRP’s ability to aggregate information across temporally distant frames. Across all retention ratios, ShaRP consistently achieves competitive performance, demonstrating its robustness and adaptability.
Cross-Backbone Generalization. To evaluate the cross-model generalization of our approach, we extend experiments to additional VLLM backbones, including LLaVA-Video [3], Qwen2-VL [4], and Qwen2.5-VL [5]. These models differ substantially in visual tokenization and feature aggregation mechanisms. For example, Qwen2-VL [4] employs 3D convolutions with MLP-based token aggregation, whereas LLaVA-Video [3] adopts newline token insertion and hierarchical pooling strategies. All methods are evaluated under identical input resolutions and comparable computational budgets. As shown in Tab. 2, we report results grouped by similar TFLOPs and retention ratios across MVBench [67, 68], LongVideoBench [69], and VideoMME [58]. ShaRP consistently achieves the best performance across the evaluated settings, demonstrating robust generalization across diverse model architectures and tokenization strategies.
| Method | Retention Ratio ↓ | TFLOPs ↓ | MVBench ↑ | LongVideo Bench ↑ | MLVU ↑ | VideoMME ↑ | Avg Score ↑ |
| VisionZip [12] | 12.8% | 4.3 | 55.6 | 51.6 | 60.9 | 55.4 | 55.9 |
[…]
| HoliTom [22] | 15%/7.5% | 4.3 | 58.4 | 55.7 | 62.1 | 57.2 | 58.4 |
[…]
| Retention Ratio ↓ | Latency (ms) | Average ↑ | |
| Prefill | Generate | ||
| LLAVA-OV | 1181.6 (1.0) | 1301.3 (1.0) | 59.0 (100%) |
| 9.31 (23.5%) | 354.3 (3.3) | 442.2 (2.9) | 58.1 (98.5%) |
| 6.03 (14.1%) | 233.7 (5.1) | 357.2 (3.6) | 57.4 (97.2%) |
| LLAVA-Video | 2404.1 (1.0) | 2796.2 (1.0) | 60.8 (100%) |
| 16.3 (23.5%) | 674.8 (3.6) | 883.5 (3.2) | 59.7 (98.2%) |
| 10.3 (14.1%) | 487.8 (4.9) | 699.3 (4.0) | 57.2 (94.1%) |
@@
Inference Efficiency. Tab. 5.2 reports the inference efficiency of ShaRP on LLaVA-OneVision [2] and LLaVA-Video [3]. ShaRP preserves strong performance while substantially reducing computation. For example, on LLaVA-OneVision, at 23.5% retention, TFLOPs drop to 9.31 and the prefill latency decreases from 1181.6 ms to 354.3 ms (3.3 speedup); at 14.1% retention, it further reduces to 233.7 ms (5.1). ShaRP also delivers up to a 4.9 prefill speedup on LLaVA-Video. To assess practical scalability, Fig. 5 reports latency as the input length increases. The near-linear growth pattern and limited incremental overhead demonstrate that ShaRP maintains efficient runtime behavior as video length increases.
@@ 5.3 Ablation Study
Ablation and Synergy Analysis. We conduct ablation experiments on our three modules—Segment-Aware Causal Masking (SegM, Sec. 4.1), Positional Bias Calibration (PosC, Sec. 4.2), and Register Token Deduplication (RegD, Sec. 4.3)—under a 23.5% token retention ratio, with results in Tab. 5. PosC alone improves the average score from 55.8 to 57.2 by enhancing focus on critical visual regions. Adding SegM (PosC+SegM) further increases performance to 57.6, mitigating attention collapse and stabilizing temporal-contextual aggregation. Combining PosC with RegD raises the score to 57.7, demonstrating RegD’s effectiveness in compressing redundant tokens and enriching structural representation. Even without PosC, SegM+RegD reaches 57.5, confirming their independent contributions. Integrating all three modules achieves the best score of 58.1, validating the complementary benefits of each component
| Method | MVBench ↑ | LongVideo Bench ↑ | MLVU ↑ | VideoMME ↑ | Avg Score ↑ |
| Original | 54.4 | 53.0 | 61.0 | 54.7 | 55.8 |
| PosC | 56.4 | 55.2 | 60.5 | 56.8 | 57.2 |
| PosC+SegM | 56.6 | 55.5 | 60.9 | 57.2 | 57.6 |
| PosC+RegD | 56.7 | 55.6 | 61.5 | 57.0 | 57.7 |
| SegM+RegD | 56.4 | 54.7 | 61.8 | 57.1 | 57.5 |
| PosC+SegM+RegD (ShaRP) | 57.1 | 55.8 | 61.9 | 57.6 | 58.1 |